A fluid undergoes a reversible adiabatic compression from pressure 1 mpa bar & volume 0.m3, according to the law pv1.3=constant. Determine the work done


1
Expert's answer
2021-02-18T18:54:06-0500

Assume that "p_1=1\ MPa, \ V_1=0.8\ m^3,\ V_2=0.4\ m^3,\ n=1.3"



"W=\frac{p_1V_1}{n-1}[1-(\frac{V_1}{V_2})^{n-1}]=\frac{1\cdot10^6\cdot 0.8}{1.3-1}[1-(\frac{0.8}{0.4})^{1.3-1}]=-172 \ (kJ)" . Answer